Your firm is planning to invest in an automated packaging plant. Harburtin Industries is an all-equity...

  1. Your firm is planning to invest in an automated packaging plant. Harburtin Industries is an all-equity firm that specializes in this business. Suppose Harburtin’s equity beta is 0.85, the risk-free rate is 4%, and the market risk premium is 5%. If your firm’s project is all equity financed, estimate its cost of capital. Consider the setting of Problem 9. You decided to look for other comparables to reduce estimation error in your cost of capital estimate. You find a second firm, Thurbinar Design, which is also engaged in a similar line of business. Thurbinar has a stock price of $20 per share, with 15 million shares outstanding. It also has $100 million in outstanding debt, with a yield on the debt of 4.5%. Thurbinar’s equity beta is 1.00.
Answer #1

NOTE: Solution provided is only for the first question as the second question is incomplete.

Harburtin Industries: All Equity - Financed, Equity Beta = B = 0.85, Risk-Free Rate = Rf = 4 % and Market Risk Premium = MRP = 5 %

As the project being considered is all-equity financed just like the firm itself, the appropriate cost of capital for the project will be the cost of equity of the firm.

Firm's cost of equity = Rf + Bx MRP = 4 + 0.85 x 5 = 8.25 %